A passenger train leaves a depot 2 hours after a freight train leaves the same depot. The passenger train is travelling 10 mph faster than the freight train. Find the rate of each train if it takes 6 hours for the freight train to catch up.
Answer by lwsshak3(11628)

A passenger train leaves a depot 2 hours after a freight train leaves the same depot. The passenger train is travelling 10 mph faster than the freight train. Find the rate of each train if it takes 6 hours for the passenger train to catch up.
***
let x=rate of freight train
x+10=rate of passenger train
distance=traveling time*rate
..
8x=6(x+10)
8x=6x+60
2x=60
x=30
x+10=40
rate of freight train=30 mph
rate of passenger train=40 mph